3. At the end of the current year, ARS Industries has the following account balances before making an adjusting entry for uncollectible accounts (10 points): Accounts Receivable has a balance $1,300,000; Allowance for Doubtful Accounts $ 7,000 Credit sales for this year $8,500,000 A. If the company uses the percent of sales method and estimates that ½ of 1% of credit sales for the year will be uncollectible, then a. Determine the bad debt expense for the period, and journalize the adjusting entries. (2 points) b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Net Realizable Value of Accounts Receivable. (3 points) Adjusted Balance Accounts Receivable _____________ Allowance for Doubtful Accounts _____________ Net Realizable Value of Accounts Receivable _____________ B. An aging of the company’s accounts receivable on December 31, and a historical analysis of the percentage of uncollectible accounts in each age class follow: Age Class Balance Percent Uncollectible Not past due $715,000 1% 1-30 days 310,000 2% 31-60 days 102,000 15% 61-90 days 76,000 30% 91-120 days 97,000 60% $1,300,000 a. Determine the bad debt expense for the period, and journalize the adjusting entries. (2 points) b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Net Realizable Value of Accounts Receivable. (3 points) Adjusted Balance Accounts Receivable _____________ Allowance for Doubtful Accounts _____________ Net Realizable Value of Accounts Receivable ____________  
Answer #1

(1)

8,500,000*1%

=85000$

85000$/2

=42500$ should be allowance for doubtful account balance at the year end

7000$ is already in balance so we would credit allowance for doubtful account more by [42500-7000]=35500$

and debit bad debt expense by that amount

journalize

date account name debit credit
December 31 bad debt expense 35500
allowance for doubtful account 35500

adjusted balances

account receivable = 1,300,000

bad debts = 35500

allowance for doubtful account = 42500

net realizable value of account = account receivable-allowance for doubtful account

=1,257,500$[1,300,000-42500]

2.

not due 715000 1% 7150
1-30 310000 2% 6200
31-60 102000 15% 15300
61-90 76000 30% 22800
91-120 97000 60% 58200
Total 1300000
allowance for doubtful account 109650

bad debt expense = 109650-7000 beginning balance

=102650

journal

Dec 31 bad debts expense 102650
allowance for doubtful account 102650

allowance for doubtful account = 109650

account receivable = 1,300,000

net account receivable = 1,300,000-109650=1190350